Grade 12th passMagical Mathematics[Interesting Approach]

The power of 3 of any index. The repetition of unit digits

Sekhar
8 Years agoGrade 12th pass

1 Answer

ARKADEEP DAS
8 Years ago
see sekhar
its nice for having law of indices
now we know 3^1=3
3^2=9
3^3=27
3^4=81
now again 3^5= 243
so we have repaeted unit’s digit after interval of 3 such that every 4th term has same unit’s place